# Spectraformer: A Unified Random Feature Framework for Transformer

### v2

@@ Abstract
Linearization of attention using various kernel approximation and kernel learning techniques has shown promise. Past methods use a subset of combinations of component functions and weight matrices within the random features paradigm. We identify the need for a systematic comparison of different combinations of weight matrix and component functions for attention learning in Transformer. In this work, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in linearized attention of the Transformer. We experiment with broad classes of component functions and weight matrices for three textual tasks in the LRA benchmark. Our experimentation with multiple combinations of component functions and weight matrices leads us to a novel combination with 23.4% faster training time and 25.2% lower memory consumption over the previous SOTA random feature Transformer, while maintaining the performance, as compared to the Original Transformer. Our code is available at: https://github.com/dukeraphaelng/spectraformer.
@@ 1 Introduction
Transformer [1] has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models. Its influence has reached beyond NLP into computer vision [2], speech processing [3] and other fields. Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism. Therefore, the performance of Transformer is tied to the effective use of attention. Attention in the Original Transformer (referred to as ‘OT’ hereafter) uses the softmax function which is quadratic in time complexity. However, softmax is not fundamental to attention. The softmax serves the role of a compatibility function, which captures the degree of compatibility between two tokens (see Equations 1 and 2). Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task [4].
@@
Replacing the softmax formulation with alternatives may allow to reduce the time complexity of attention computation. In this paper, we focus on the subset of alternatives that employs kernel functions. Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters [5]. [4] show that attention can be formulated using kernels and then be linearized [6], given the feature map of the respective attention kernel. Random feature-based algorithms [7] are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel. A random feature consists of three components: component function, weight matrix, and component weight (see Section 3). Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by [8]. This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance [9, 8, 10, 11]; (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded [8, 12, 13]; (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation [14]. (A) and (B) seek to improve approximation quality via approximation error and variance reduction. (C) dispenses with the approximation of the attention kernel (typically the softmax), and makes the kernel itself learnable, to attain a better performance. However, the three have been explored separately, creating a situation where there are different ways to construct the random features. Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature. Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer [9, 8, 10, 11], ignoring those which are popular in the kernel methods literature (e.g., [7]). These limitations necessitate a unified framework that allows the comparison of different combinations in search of the best-performing one.
Figure 1: Peak memory consumption (x-axis), average accuracy (y-axis) on three textual LRA [15] tasks: ListOps, Text, Retrieval. Color indicates training time, the darker the faster, the brighter the slower (see legend). Models compared are the best Spectraformer variant (PosRF-MM) (boldened), the OT, and other efficient Transformers (RF and non-RF) in the literature. Values for non-RF Transformers are taken from [16].
@@
We introduce Spectraformer, a unified framework of random feature-based attention in Transformer. It utilizes spectral analysis-inspired random features to construct linearized attention in Transformers. Spectraformer allows for the experimentation of various combinations of weight matrices and component functions. We will open-source our codebase upon acceptance. Through our benchmarking experiments on three textual Long Range Arena [15] tasks and 18 combinations of weight matrices and component functions, Spectraformer is able to find novel combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption. Figure 1 compares our best-known Spectraformer combination (PosRF-MM) against the previous SOTA models in random feature Transformer, efficient Transformers, and the OT. The x and y axes denote peak memory and average accuracy, while the colour of the circle indicates the memory consumption. Spectraformer (specifically, PosRF-MM) is more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# [13]), OPRF-ORF (a.k.a., FAVOR++ [12]), PosRF-ORF (a.k.a., FAVOR+ [8]) and maintains a comparable accuracy against the OT, whilst significantly reducing training time and memory consumption over both. However, it underperforms in accuracy and training time against Skyformer [16] and Nyströmformer [17]. We hope that Spectraformer will enable the mechanism to benchmark more random feature-based Transformers for these and other tasks. Our contributions are:
@@
    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark. Our best model has comparable accuracy to the OT whilst significantly reducing training time and peak memory consumption.
@@
    This is the first work to compare random feature Transformers with other efficient Transformers.
@@ 2.1 Kernelized Attention
OT attention [1] can be defined using the unified attention model [18] as follows. Given a collection of inputs and a learning objective, attention mechanism learns the attention matrix A which captures the associative information between pairs of tokens in a sequence of length $N$, with each row $\textbf{A}_{i}$ being the representation of a token $i$ with every other token $j\in[1,N]$. This is done by creating three learnable representations of each token $i$ called query ($\boldsymbol{q}_{i}$), key ($\boldsymbol{k}_{i}$), and value ($\boldsymbol{v}_{i}$). The attention representation $\textbf{A}_{i}$ is calculated as a weighted sum of value $\boldsymbol{v}_{j}$ given attention weights $a_{j}$. The general equation for $\textbf{A}_{i}$ is given as:
$\begin{array}[]{cccc}\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})}{\sum_{l}f(\boldsymbol{q}_{i},\boldsymbol{k}_{l})}\boldsymbol{v}_{j}=\sum_{j=1}^{N}a_{j}\boldsymbol{v}_{j};&a=g(e);&e=f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})\end{array}$ (1)
Attention weights are generated using a distribution function $g$ applied on an energy score $e$. The energy score is captured by applying a compatibility function on a query $\boldsymbol{q}_{i}$ and a set of keys K, in order to compute a score of a specific relationship between the pairs. The attention in the OT is called the ‘scaled dot-product’ which defines the compatibility function $f$ and the distribution function $g$ as follows:
$\begin{array}[]{cc}f(\boldsymbol{q}_{i},\textbf{K})=\frac{\boldsymbol{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\boldsymbol{x}_{j})=\frac{\exp(\boldsymbol{x}_{j})}{\sum_{l}\exp(\boldsymbol{x}_{l})}\end{array}$ (2)
@@
$\begin{array}[]{cc}\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\boldsymbol{q}_{i}\boldsymbol{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\boldsymbol{q}_{i}\boldsymbol{k}_{l}^{T}/\sqrt{d_{k}})}\boldsymbol{v}_{j}&A=softmax(\frac{\boldsymbol{QK}^{T}}{\sqrt{d_{k}}})\boldsymbol{V}\end{array}$ (3)
with $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}$ being the matrix of queries, keys, and values respectively with each row being $\boldsymbol{q}_{i},\boldsymbol{k}_{i},\boldsymbol{v}_{i}$.
We now discuss the intrinsic connection between kernel and attention. With the compatibility function $f$ as a kernel $f(\boldsymbol{q}_{i},\boldsymbol{k}_{j})=\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})$, Equation 1 can be rewritten as:
$\boldsymbol{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})}{\sum_{l}\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l})}\boldsymbol{v}_{j}$ (4)
As shown in [4], this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\boldsymbol{k}_{j}|\boldsymbol{q}_{i})}[\boldsymbol{v}_{j}|X=\boldsymbol{k}_{j}]=\boldsymbol{A}_{i}$, $l_{i}\left(\boldsymbol{k}_{j})=p(\boldsymbol{k}_{j}|\boldsymbol{q}_{i}\right)$, $Y_{i}=\boldsymbol{v}_{j}$:
$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\boldsymbol{u}]=\sum_{i=1}^{N}l_{i}(\boldsymbol{u})Y_{i};&l_{i}(\boldsymbol{u})=\frac{\mathcal{K}(\frac{\boldsymbol{u}-\boldsymbol{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\boldsymbol{u}-\boldsymbol{x}_{j}}{h})}\end{array}$ (5)
@@
$\mathcal{K}_{softmax}(\boldsymbol{q}_{i},\boldsymbol{k}_{j})=\exp(\boldsymbol{q}_{i}^{T}\boldsymbol{k}_{j}/\sqrt{d_{k}})$ (6)
@@ 2.2 Linearizing Kernelized Attention via Random Features
Since a kernel is the inner product of two vectors in some space with respect to some feature map $\phi$, i.e., $\mathcal{K}(\boldsymbol{x},\boldsymbol{y})=\phi(\boldsymbol{x})\phi(\boldsymbol{y})^{T}$, then we can rewrite Equation 4 as:
$\begin{array}[]{c}\boldsymbol{A}_{i}=\frac{\sum_{j}\phi(\boldsymbol{q}_{i})\phi(\boldsymbol{k}_{j})^{T}\boldsymbol{v}_{j}}{\sum_{l}\phi(\boldsymbol{q}_{i})\phi(\boldsymbol{k}_{l})^{T}}=\frac{\phi(\boldsymbol{q}_{i})\sum_{j}\phi(\boldsymbol{k}_{j})\otimes\boldsymbol{v}_{j}}{\phi(\boldsymbol{q}_{i})\sum_{l}\phi(\boldsymbol{k}_{l})^{T}}\end{array}$ (7)
If we pre-compute $\sum_{j}\phi(\boldsymbol{k}_{j})\otimes\boldsymbol{v}_{j}$ and $\sum_{l}\phi(\boldsymbol{k}_{l})^{T}$, the entire term becomes $O(1)$ and we only need to compute Equation 7 $N$ times for each $i$, resulting in an attention matrix calculated in linear time $O(N)$. We now refer to Equation 7 as linearized attention [6].
All that is left is to retrieve the feature map $\phi$, the process of which is called kernel approximation (see Section 2.3), which corresponds to the kernels we specify. There are many popular kernels from which we can choose, but typically, we want to approximate the softmax kernel in Equation 6. However, since the softmax relates to the $\mathcal{K}_{RBF}(\boldsymbol{x},\boldsymbol{y})=\exp(-\frac{||\boldsymbol{x}-\boldsymbol{y}||^{2}}{2})$ via the following equation:
$\mathcal{K}_{softmax}(\boldsymbol{x},\boldsymbol{y})=\exp(\frac{||\boldsymbol{x}||^{2}}{2})\mathcal{K}_{RBF}(\boldsymbol{x},\boldsymbol{y})\exp(\frac{||\boldsymbol{y}||^{2}}{2})$ (8)
@@
Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in [4]. Then kernel learning would be a more suitable option, the detail of which is discussed in Section 2.4. Notwithstanding, we first introduce the main kernel approximation technique: random features.

### v3

@@ Abstract
Linearization of attention using various kernel approximation and kernel learning techniques has shown promise. Past methods use a subset of combinations of component functions and weight matrices within the random features paradigm. We identify the need for a systematic comparison of different combinations of weight matrices and component functions for attention learning in Transformer. In this work, we introduce Spectraformer, a unified framework for approximating and learning the kernel function in linearized attention of the Transformer. We experiment with broad classes of component functions and weight matrices for three textual tasks in the LRA benchmark. Our empirical findings indicate that different kernels are good at different tasks and that kernel choice is fundamental to performant models. Our code is available at: https://anonymous.4open.science/r/spectraformer-8A97.
@@ 1 Introduction
Transformer (Vaswani et al., 2017) has revolutionized the landscape of natural language processing (NLP) and forms the basis of almost all state-of-the-art language models. Its influence has reached beyond NLP into computer vision (Han et al., 2023), speech processing (Lin et al., 2022) and other fields. Compared to its precursors like the LSTMs, Transformer leverages parallelism since it is fully based on the attention mechanism. Therefore, the performance of Transformer is tied to the effective use of attention. Attention in the Original Transformer (referred to as ‘OT’ hereafter) uses the softmax function which is quadratic in time complexity. However, softmax is not fundamental to attention. The softmax serves the role of a compatibility function, which captures the degree of compatibility between two tokens (see Equations 1 and 2). Extensive experiments with other compatibility functions show that the softmax is not the only possibility, and that the ‘best’ compatibility function depends on the task (Tsai et al., 2019).
@@
Replacing the softmax formulation with alternatives may allow us to reduce the time complexity of attention computation. In this paper, we focus on the subset of alternatives that employs kernel functions. Kernel functions have been historically used in machine learning algorithms to simplify the estimation of parameters (Hofmann et al., 2008). Tsai et al. (2019) show that attention can be formulated using kernels and then be linearized (Katharopoulos et al., 2020), given the feature map of the respective attention kernel. Random feature-based algorithms (Liu et al., 2022) are a family of algorithms, inspired by spectral analysis, which provides the feature map associated with a kernel. A random feature consists of three components: component function, weight matrix, and component weight (see Section 3). Due to their robustness and flexibility, random features give rise to ‘random feature Transformers’, the family of Transformers with linearized attention via random features (hereby abbreviated as RF), first with the work by Choromanski et al. (2021). This is seen to lead to three strands of research: (A) Optimization of weight matrices either by making the matrices have lower time or space complexity, or by incorporating beneficial properties like orthogonality to reduce variance (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023); (B) Enhancement of component functions by either: (i) finding a function with a tighter variance or approximation bound, or (ii) be engineered to output with numerical stability and be bounded (Choromanski et al., 2021; Likhosherstov et al., 2022; 2023); (C) Parameterization of weight matrices to perform kernel learning instead of kernel approximation (Chowdhury et al., 2022). (A) and (B) seek to improve approximation quality via approximation error and variance reduction. (C) dispenses with the approximation of the attention kernel (typically the softmax), and makes the kernel itself learnable, to attain a better performance. However, the three have been explored separately, creating a situation where there are different ways to construct the random features. Additionally, past works only explore certain combinations of component functions and weight matrices, creating disjoint overlaps and gaps in the literature. Finally, the weight matrices are often those explicitly designed in the context of kernelized attention in the Transformer (Yu et al., 2016; Choromanski et al., 2021; 2022; Reid et al., 2023), ignoring those which are popular in the kernel methods literature (e.g., Liu et al. (2022)). These limitations necessitate a unified framework that allows the comparison of different combinations in search of the best-performing one.
@@
We introduce Spectraformer, a unified framework of random feature-based attention in Transformer. It utilizes spectral analysis-inspired random features to construct linearized attention in Transformers. Spectraformer allows for the experimentation of various combinations of weight matrices and component functions. We will open-source our codebase upon acceptance. Through our benchmarking experiments on three textual Long Range Arena (Tay et al., 2021) tasks and 18 combinations of weight matrices and component functions, Spectraformer enables novel combinations of component functions and weight matrices from the kernel literature to be experimented easily in the Transformer setting. This process makes possible the discovery of combinations that outperform existing random feature-based Transformers in terms of performance, training time, and peak memory consumption. Various novel Spectraformer combinations are more accurate than the previous SOTA models, SADERF-ORF (a.k.a., FAVOR# (Likhosherstov et al., 2023)), OPRF-ORF (a.k.a., FAVOR++ (Likhosherstov et al., 2022)), PosRF-ORF (a.k.a., FAVOR+ (Choromanski et al., 2021)), whilst significantly reducing training time and memory consumption over both. We hope that Spectraformer will enable the mechanism to benchmark more random feature-based Transformers for these and other tasks. Our contributions are:
@@
    We demonstrate how Spectraformer can discover novel SOTA models in the random feature Transformer family for the LRA benchmark and that these models are more performant than existing random feature Transformers.
@@
    We empirically identify the relationship between performant combinations and task characteristics, aiding the right kernel choice when tackling a task.
  4. 4.
    Our code is currently available for review, but they will be publicly released upon acceptance. We detail guide on how to incorporate future kernel works into Spectraformer. With this work, every future kernel work that is introduced can easily be incorporated into the transformer.
@@ 2 Preliminaries
We first cover the theoretical foundations to this work. A list of commonly used acronyms is available in Table 1.
Table 1: Table of acronyms
Acronym Description
OT (Vaswani et al., 2017) Original Transformer
RF (Rahimi & Recht, 2007) Random Feature method
RBF Radial Basis Function kernel (also called the Gaussian kernel)
TrigRF (Rahimi & Recht, 2007) Trigonometric RF (component function)
PosRF (Choromanski et al., 2021) Positive RF (component function)
OPRF (Likhosherstov et al., 2022) Optimized positive RF (component function)
SADERF (Likhosherstov et al., 2023) Simplified Asymmetric Dense-Exponential RF (component function)
ORF (Yu et al., 2016) Orthogonal RF (weight matrix)
SORF (Yu et al., 2016) Structural ORF (weight matrix)
QMC (Avron et al., 2016) Quasi- Monte Carlo (weight matrix)
MM (Shen et al., 2017) (Liu et al., 2022) Moment Matching (weight matrix)
SGQ (Dao et al., 2017) Sparse Grid Quadrature (weight matrix)
FastFoodL(Yang et al., 2015) FastFood Learnable (weight matrix)
@@ 2.1 Kernelized Attention
OT attention (Vaswani et al., 2017) can be defined using the unified attention model (Galassi et al., 2021) as follows. Given a collection of inputs and a learning objective, attention mechanism learns the attention matrix A which captures the associative information between pairs of tokens in a sequence of length $N$, with each row $\textbf{A}_{i}$ being the representation of a token $i$ with every other token $j\in[1,N]$. This is done by creating three learnable representations of each token $i$ called query ($\bm{q}_{i}$), key ($\bm{k}_{i}$), and value ($\bm{v}_{i}$). The attention representation $\textbf{A}_{i}$ is calculated as a weighted sum of value $\bm{v}_{j}$ given attention weights $a_{j}$. The general equation for $\textbf{A}_{i}$ is given as:
$\begin{array}[]{cccc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{f(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}f(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}=\sum_{j=1}^{N}a_{j}\bm{v}_{j};&a=g(e);&e=f(\bm{q}_{i},\bm{k}_{j})\end{array}$ (1)
Attention weights are generated using a distribution function $g$ applied on an energy score $e$. The energy score is captured by applying a compatibility function on a query $\bm{q}_{i}$ and a set of keys K, in order to compute a score of a specific relationship between the pairs. The attention in the OT is called the ‘scaled dot-product’ which defines the compatibility function $f$ and the distribution function $g$ as follows:
$\begin{array}[]{cc}f(\bm{q}_{i},\textbf{K})=\frac{\bm{q}_{i}^{T}\textbf{K}}{\sqrt{d_{k}}}&g(\bm{x}_{j})=\frac{\exp(\bm{x}_{j})}{\sum_{l}\exp(\bm{x}_{l})}\end{array}$ (2)
@@
$\begin{array}[]{cc}\bm{A}_{i}=\sum_{j=1}^{N}\frac{\exp(\bm{q}_{i}\bm{k}_{j}^{T}/\sqrt{d_{k}})}{\sum_{l}\exp(\bm{q}_{i}\bm{k}_{l}^{T}/\sqrt{d_{k}})}\bm{v}_{j}&A=softmax(\frac{\bm{QK}^{T}}{\sqrt{d_{k}}})\bm{V}\end{array}$ (3)
with $\bm{Q},\bm{K},\bm{V}$ being the matrix of queries, keys, and values respectively with each row being $\bm{q}_{i},\bm{k}_{i},\bm{v}_{i}$.
We now discuss the intrinsic connection between kernel and attention. With the compatibility function $f$ as a kernel $f(\bm{q}_{i},\bm{k}_{j})=\mathcal{K}(\bm{q}_{i},\bm{k}_{j})$, Equation 1 can be rewritten as:
$\bm{A}_{i}=\sum_{j=1}^{N}\frac{\mathcal{K}(\bm{q}_{i},\bm{k}_{j})}{\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l})}\bm{v}_{j}$ (4)
As shown in Tsai et al. (2019), this is indeed the Nadaraya-Watson Kernel Estimator, with $\mathbb{E}_{p(\bm{k}_{j}|\bm{q}_{i})}[\bm{v}_{j}|X=\bm{k}_{j}]=\bm{A}_{i}$, $l_{i}\left(\bm{k}_{j})=p(\bm{k}_{j}|\bm{q}_{i}\right)$, $Y_{i}=\bm{v}_{j}$:
$\begin{array}[]{ccc}\mathbb{E}_{l_{i}}[Y_{i}|X=\bm{u}]=\sum_{i=1}^{N}l_{i}(\bm{u})Y_{i};&l_{i}(\bm{u})=\frac{\mathcal{K}(\frac{\bm{u}-\bm{x}_{i}}{h})}{\sum_{j=1}^{n}\mathcal{K}(\frac{\bm{u}-\bm{x}_{j}}{h})}\end{array}$ (5)
@@
$\mathcal{K}_{softmax}(\bm{q}_{i},\bm{k}_{j})=\exp(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{k}})$ (6)
@@ 2.2 Linearizing Kernelized Attention via Random Features
Since a kernel is the inner product of two vectors in some space with respect to some feature map $\phi$, i.e., $\mathcal{K}(\bm{x},\bm{y})=\phi(\bm{x})\phi(\bm{y})^{T}$, then we can rewrite Equation 4 as:
$\begin{array}[]{c}\bm{A}_{i}=\frac{\sum_{j}\phi(\bm{q}_{i})\phi(\bm{k}_{j})^{T}\bm{v}_{j}}{\sum_{l}\phi(\bm{q}_{i})\phi(\bm{k}_{l})^{T}}=\frac{\phi(\bm{q}_{i})\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}}{\phi(\bm{q}_{i})\sum_{l}\phi(\bm{k}_{l})^{T}}\end{array}$ (7)
If we pre-compute $\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}$ and $\sum_{l}\phi(\bm{k}_{l})^{T}$, the entire term becomes $O(1)$ and we only need to compute Equation 7 $N$ times for each $i$, resulting in an attention matrix calculated in linear time $O(N)$. We now refer to Equation 7 as linearized attention (Katharopoulos et al., 2020).
All that is left is to retrieve the feature map $\phi$, the process of which is called kernel approximation (see Section 2.3), which corresponds to the kernels we specify. There are many popular kernels from which we can choose, but typically, we want to approximate the softmax kernel in Equation 6. However, since the softmax relates to the $\mathcal{K}_{RBF}(\bm{x},\bm{y})=\exp(-\frac{||\bm{x}-\bm{y}||^{2}}{2})$ via the following equation:
$\mathcal{K}_{softmax}(\bm{x},\bm{y})=\exp(\frac{||\bm{x}||^{2}}{2})\mathcal{K}_{RBF}(\bm{x},\bm{y})\exp(\frac{||\bm{y}||^{2}}{2})$ (8)
@@
Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in Tsai et al. (2019). Then kernel learning would be a more suitable option, the detail of which is discussed in Section 2.4. Notwithstanding, we first introduce the main kernel approximation technique: random features.